OpenStudy (anonymous):

determine the centre and radius of the circle with equation 3x^2+3y^2-5x+6y+3=0 [using completion of squares] please show all working...thank you :)

OpenStudy (anonymous):

this is the same crap im doing lol good luck i have got no help

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

This is the first part

OpenStudy (amistre64):

(3x^2 -5x) + (3y^2 +6y) +3 = 0 3(x^2 -5/3 x) + 3(y^2 +2y) +3 = 0 3(x^2 -5/3x +(5/6)^2) + 3(y^2 +2y +1) +3-3((5)/6)^2-3(1) = 0 3(x -5/6)^2 + 3(y+1)^2 -3(25/36) = 0 3(x -5/6)^2 + 3(y+1)^2 -25/12 = 0 perhaps?

OpenStudy (anonymous):

That is the second part

OpenStudy (amistre64):

3(x -5/6)^2 + 3(y+1)^2 -25/12 = 0 ; \3 (x -5/6)^2 + (y+1)^2 -25/36 = 0 (x -5/6)^2 + (y+1)^2 = 25/36 maybe :)

OpenStudy (amistre64):

center: (5/6 , -1) radius = 5/6

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

thanks emunrradtvamg and amistre but i dont see how you guys arrived at your final answer because of the genral from of the equation i saw how we got the radius because its r but k is the center?

OpenStudy (anonymous):

Look at the second picture I send you. Im gonna try to explain

OpenStudy (anonymous):

i look at it i saw the general form of the equation (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Exactly As you can see, inside of the first term you have (x-5/6)^2 that is equivalent to (x-h)^2. So you can realize from that that h = 5/6. (y+1) is equivalent to (y-k)^2. So k = -1 (look carefully at the formula) the center is (h,k) = (5/6,-1)

OpenStudy (anonymous):

oh okay i get it now thank you so much emunrradtvamg and amistre!!!